# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

### v1

@@ IV-A Datasets and Implementation Details
DRSformer [204] CVPR’23 33M 29.02/0.968 35.89/0.970 33.28/0.921 30.55/0.862 27.58/0.786 31.26/0.902 Specific
@@
PromptIR [65] NeurIPS’23 33M 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 prompt for AiOIR
TextPromptIR [94] arXiv’23 - 31.65/0.978 38.41/0.982 34.17/0.936 31.52/0.893 28.26/0.805 32.80/0.919 textual prompt for AiOIR
@@
NDR [96] arXiv’23 28M 28.64/0.962 35.42/0.969 34.01/0.932 31.36/0.887 28.10/0.798 31.51/0.910 neural degradation representation
DA-CLIP [76] ICLR’24 125M 29.46/0.963 36.28/0.968 30.02/0.821 24.86/0.585 22.29/0.476 -/- degradation-aware vision-language model
@@
$\text{Art}_{\text{AirNet}}$ [135] ACM MM’24 9M 30.56/0.977 37.74/0.981 34.02/0.934 31.37/0.890 28.12/0.802 32.36/0.917 via multi-task collaboration
$\text{Art}_{\text{PromptIR}}$ [135] ACM MM’24 33M 30.83/0.979 37.94/0.982 34.06/0.934 31.42/0.891 28.14/0.801 32.49/0.917 via multi-task collaboration
AdaIR [78] arXiv’24 29M 31.06/0.980 38.64/0.983 34.12/0.935 31.45/0.892 28.19/0.802 32.69/0.918 frequency mining and modulation
CAPTNet [95] TCSVT’24 - 29.28/- 37.86/- -/- 30.75/- -/- -/- prompt-based&ingredient-Oriented
InstructIR-3D [62] ECCV’24 16M 30.22/0.959 37.98/0.978 34.15/0.933 31.52/0.890 28.30/0.804 32.43/0.913 natural language prompts
InstructIR-5D [62] ECCV’24 16M 27.10/0.956 36.84/0.973 34.00/0.931 31.40/0.887 28.15/0.798 31.50/0.909 natural language prompts
@@
Perceive-IR [111] arXiv’24 42M 30.87/0.975 38.29/0.980 34.13/0.934 31.53/0.890 28.31/0.804 32.63/0.917 quality-aware degradation
UniProcessor [210] arXiv’24 - 31.66/0.979 38.17/0.982 34.08/0.935 31.42/0.891 28.17/0.803 32.70/0.918 support multimodal control
@@ IV-B Experiments and Analysis
To demonstrate the superiority of different AiOIR models, we provide an objective quality comparison in Tab. III, Tab. IV, Tab.V. The evaluation metrics are composed of PSNR, SSIM [36]. Specifically, we summarize the experimental results of different methods under three common experimental settings in the field of AiOIR, i.e., Setting1 for three distinct tasks in Tab. III: dehazing, deraining, denoising (the noise level $\sigma=15$, $\sigma=25$, $\sigma=50$), Setting2 for five distinct tasks in Tab. IV: dehazing, deraining, denoising (the noise level $\sigma=25$), deblurring, and low-light enhancement, Setting3 for synthetic deweathering (the All-weather datasets [64]) in Tab. V: snow, rain+fog, raindrop, Setting4 for real-world deweathering (the WeatherStream dataset [173]) in Tab. VI: haze, rain, snow. To compare the computation cost and network complexity, we also measure the parameters for partial methods. The results align closely with the original paper, and for settings not tested in the original, we choose results with higher evaluation metrics.
@@
For Setting1 in Tab. III, PIP [72] and TextPromptIR [94] achieve the best average performance, indicating strong generalization across different degradation types. PIP achieves PSNR/SSIM of 32.91/0.920, while TextPromptIR closely follows with 32.80/0.919, which reflects the effectiveness of prompt-based strategies and multimodal prompt. For dehazing, Instruct-IPT [116] achieves the highest PSNR of 39.95, significantly outperforming others, followed by MEASNet [110] and DyNet [185]. However, MEASNet and DyNet stand out with the highest PSNR values of 39.00 and 38.71 for deraining, respectively. Many recent models adopt complex mechanisms such as frequency-aware transformations [209], multi-expert selection [110], and hypernetwork-based architectures [189], suggesting a growing emphasis on specialized designs to tackle varying degradation patterns.
@@
For Setting2 in Tab. IV, $\text{PIP}_{\text{Restormer}}$ also achieves the best average performance across tasks, particularly excelling in dehazing and deraining. MEASNet and DaAIR [187] also show strong results, with high average scores, indicating their robustness across multiple degradation types. Prompt-based models like $\text{PIP}_{\text{Restormer}}$ and PromptIR [65] are among the top performers, suggesting that prompt learning strategies and AiOIR are becoming dominant approaches for tackling multiple degradations simultaneously. Degradation-awareness seems to be a key factor, with models like DaAIR and Perceive-IR [111] showing that tailoring the restoration process based on the specific type of degradation improves overall performance. The balance between model size (params) and performance varies significantly. For example, smaller models like TAPE [59] perform reasonably well in certain tasks, indicating efficient design, while larger models like Gridformer [104] leverage more complex architectures for more consistent results across tasks.
@@
TABLE IV: Performance Comparisons of AiOIR Models on Five Challenging Datasets.
Denoising results are reported for the noise level $\sigma=25$.
@@
  Method Venue Params. Dehazing Deraining Denoising Deblurring Low-light Average Approach
@@
$\text{PIP}_{\text{NAFNet}}$ [72] arXiv’23 - 31.75/0.978 37.67/0.980 31.25/0.878 28.08/0.853 23.37/0.854 30.66/0.899 prompt-in-prompt learning
@@
PromptIR [65] NeurIPS’23 33M 26.54/0.949 36.37/0.970 31.47/0.886 28.71/0.881 22.68/0.832 29.15/0.904 prompt for AiOIR
@@
$\text{DASL}_{\text{DGINet}}$ [215] arXiv’23 17M 25.33/0.943 36.96/0.972 31.23/0.885 27.23/0.836 21.78/0.824 28.51/0.892 decomposition ascribed synergistic
$\text{DASL}_{\text{MPRNet}}$ [215] arXiv’23 5M 23.64/0.924 34.93/0.961 30.99/0.883 26.04/0.788 20.06/0.805 27.13/0.872 decomposition ascribed synergistic
@@
For Setting3 in Tab. V, the AWRCP model [68] achieves the best results across all three weather conditions, demonstrating codebook priors’ strong capacity to handle complex weather degradations. Models like WeatherDiff [69] and TKMANet [100] also perform well, particularly in handling snow and foggy conditions. Finally, older methods like All-in-One [56] perform significantly lower compared to more recent models like AWRCP and Transweather [64], with an average PSNR of only 28.05 dB, showing that recent advancements.
For Setting4 in Tab. VI, despite Transweather’s larger parameter count, struggles to deliver competitive performance. In contrast, though more parameter-efficient, AirNet still falls short in overall restoration quality. TKMANet and WGWS-Net show a notable improvement in both metrics, balancing effectiveness and efficiency, with WGWS-Net standing out as a lightweight model that achieves higher restoration quality. The best performance is observed with the model by Yang et al. [74], which leverages language-driven techniques to achieve superior results across all weather conditions. Overall, these results demonstrate a clear progression in multi-weather restoration, suggesting that the field is advancing towards more sophisticated and robust techniques across diverse conditions.
TABLE V: Performance Comparisons of AiOIR Models on All-Weather Datasets [64]. Below the dividing line in the table is AiOIR methods.
@@
TABLE VI: Performance Comparisons of AiOIR Models on WeatherStream Datasets [173]. Below the dividing line in the table is AiOIR methods.
@@
LDR [74] 23.11/0.809 24.42/0.818 23.12/0.838 23.55/0.822 -

### v2

@@ IV-A Datasets and Evaluation Metrics
DRSformer [19] CVPR’23 34M 29.02/0.968 35.89/0.970 33.28/0.921 30.55/0.862 27.58/0.786 31.26/0.902 Specific
@@
PromptIR [58] NeurIPS’23 36M 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 prompt for AiOIR
@@
NDR [62] TIP’24 28M 28.64/0.962 35.42/0.969 34.01/0.932 31.36/0.887 28.10/0.798 31.51/0.910 neural degradation representation
$\text{Art}_{\text{PromptIR}}$ [130] ACM MM’24 36M 30.83/0.979 37.94/0.982 34.06/0.934 31.42/0.891 28.14/0.801 32.49/0.917 via multi-task collaboration
@@
LoRA-IR [155] arXiv’24 - 30.68/0.961 37.75/0.979 34.06/0.935 31.42/0.891 28.18/0.803 32.42/0.914 mixture of low-rank experts
Up-Restorer [166] AAAI’25 28M 30.68/0.977 36.74/0.978 33.99/0.933 31.33/0.888 28.07/0.799 32.16/0.915 ADMM-solver-based design
$\text{TUR}_{\text{PromptIR}}$ [131] AAAI’25 36M 31.17/0.978 38.57/0.984 34.06/0.932 31.40/0.887 28.13/0.797 32.67/0.916 a novel loss function
AdaIR [79] ICLR’25 29M 31.06/0.980 38.64/0.983 34.12/0.935 31.45/0.892 28.19/0.802 32.69/0.918 frequency mining and modulation
DA-RCOT [233] TPAMI’25 50M 31.26/0.977 38.36/0.983 33.98/0.934 31.33/0.890 28.10/0.801 32.60/0.917 as an optimal transport problem
ABAIR [234] arXiv’24 41M 33.53/0.984 38.69/0.982 34.18/0.935 31.38/0.890 28.25/0.804 33.21/0.919 a three-phase approach
MoCE-IR [153] CVPR’25 25M 31.34/0.979 38.57/0.984 34.11/0.932 31.45/0.888 28.18/0.800 32.73/0.917 mixture-of-complexity-experts framework
MTAIR [235] arXiv’24 - 31.34/0.983 39.15/0.984 34.14/0.936 31.50/0.893 28.24/0.805 32.87/0.920 Mamba-Transformer cross-dimensional
$\text{DCPT}_{\text{PromptIR}}$ [236] ICLR’25 35M 31.91/0.981 38.43/0.983 34.17/0.933 31.53/0.889 28.30/0.802 32.87/0.918 learn to classify degradation
RamIR [237] APPL INTELL’25 22M 31.29/0.977 38.16/0.981 34.04/0.931 31.61/0.891 28.19/0.801 32.65/0.916 prompt-driven Mamba-based
Cat-AIR [203] arXiv’25 - 31.49/0.980 38.43/0.983 34.11/0.935 31.44/0.892 28.14/0.803 32.72/0.919 content and task-aware framework
DSwinIR [238] arXiv’25 24M 31.86/0.980 37.73/0.983 34.12/0.933 31.59/0.890 28.31/0.803 32.72/0.917 deformable sliding window transformer
$\text{CPL}_{\text{PromptIR}}$ [116] arXiv’25 36M 31.27/0.980 38.77/0.985 34.15/0.933 31.50/0.889 28.23/0.800 32.78/0.917 contrastive prompt regularization
AnyIR [239] arXiv’25 6M 31.75/0.981 38.61/0.984 34.12/0.936 31.46/0.893 28.20/0.804 32.83/0.920 local-global gated mechanism
MIRAGE [240] arXiv’25 10M 31.86/0.981 38.94/0.985 34.12/0.935 31.46/0.891 28.19/0.803 32.91/0.919 within the SPD manifold space
BaryIR [241] arXiv’25 - 31.33/0.980 38.95/0.984 34.16/0.935 31.54/0.892 28.25/0.802 32.85/0.919 in continuous barycenter space
$\text{SIPL}_{\text{PromptIR}}$ [242] arXiv’25 39M 31.09/0.977 38.43/0.984 34.12/0.933 31.48/0.889 28.22/0.800 32.67/0.917 privilege learning
[…]
Perceive-IR [84] $\bigstar$ arXiv’24 42M+86M 30.87/0.975 38.29/0.980 34.13/0.934 31.53/0.890 28.31/0.804 32.63/0.917 quality-aware degradation
VLU-Net [149] $\bigstar$ CVPR’25 35M+88M 30.71/0.980 38.93/0.984 34.13/0.935 31.48/0.892 28.23/0.804 32.70/0.919 vision-language gradient descent-driven
[…]
Note: The dash (“-”) in the “Params” column indicates that there is no corresponding source code. $\bigstar$ denotes models assisted by foundation models (e.g., CLIP, BERT, or DINO); $\bigoplus$ denotes models assisted by additional modalities (e.g., text or depth). These notations are consistent across all subsequent tables.
@@ IV-B Experiments and Analysis
To demonstrate the superiority of different AiOIR models, we provide an objective quality comparison in Tab. II, Tab. III, Tab.IV, Tab.V, Tab.VI. The evaluation metrics are composed of PSNR, SSIM [212]. Specifically, we summarize the experimental results of different methods under five common experimental settings in the field of AiOIR, i.e., Setting1 for three-degradation scenarios in Tab. II: dehazing, deraining, denoising (the noise level $\sigma=15$, $\sigma=25$, $\sigma=50$), Setting2 for five-degradation scenarios in Tab. III: dehazing, deraining, denoising (the noise level $\sigma=25$), deblurring, and low-light enhancement, Setting3 for synthetic weather-related scenarios (the All-Weather dataset [29]) in Tab. IV: snow, rain+fog and raindrop, Setting4 for real-world weather-related scenarios in Tab. V: haze, rain and snow, and Setting5 for mixed-degradation scenarios (the CDD-11 dataset [31]) in Tab. VI: low-light, haze, rain, snow, low+haze, low+rain, low+snow, haze+rain, haze+snow, low+haze+rain, and low+haze+snow. The results align closely with the original paper.
@@
For Setting1 in Tab. II, ABAIR [234] and DFPIR [243] achieve the highest average performance both without and with the assistance of foundational models, respectively, indicating strong generalization across diverse degradations. ABAIR attains an average PSNR/SSIM of 32.91/0.920, while DFPIR follows closely with 32.88/0.919, underscoring that careful model design and training strategies are as crucial as leveraging VLMs. For dehazing, ABAIR also achieves the highest PSNR of 33.53, significantly outperforming others, followed by DaAIR [230] at 32.30 and PIP [89] at 32.09. MTAIR [235] and MEASNet [82] stand out with the highest PSNR values of 39.15 and 39.00 for deraining, respectively. Many recent models adopt complex mechanisms such as frequency-aware transformations [232], optimal transport [233], deep unfolding networks [149], and hypernetwork-based architectures [205], suggesting a growing emphasis on specialized designs to tackle varying degradation patterns.
@@
For Setting2 in Tab. III, D3Net [244] achieves the best average performance across tasks, particularly excelling in dehazing and low-light enhancement. ABAIR and DFPIR also show strong results, with high average scores, indicating their robustness across multiple degradation types. Prompt-based models such as $\text{CPL}_{\text{PromptIR}}$ [116] and CyclicPrompt [148] are among the top performers, suggesting that prompt learning strategies remain a dominant paradigm for tackling multiple degradations simultaneously. Degradation-awareness remains crucial, with models like DaAIR and Perceive-IR [84] tailoring the restoration process to specific degradation characteristics and yielding excellent performance. The trade-off between model size and performance is also evident. Lightweight designs such as TAPE [245] achieve competitive results in certain tasks, whereas more complex architectures like Gridformer [67] deliver more reliable performance across all tasks.
@@
TABLE III: Performance Comparisons of AiOIR Models on Five Challenging Datasets. Denoising results are reported for the noise level $\sigma=25$.
@@
  Method Venue & Year Params Dehazing Deraining Denoising Deblurring Low-light Average Approach
@@
PromptIR [58] NeurIPS’23 36M 26.54/0.949 36.37/0.970 31.47/0.886 28.71/0.881 22.68/0.832 29.15/0.904 prompt for AiOIR
@@
$\text{Art}_{\text{PromptIR}}$ [130] ACM MM’24 36M 29.93/0.908 22.09/0.891 29.43/0.843 25.61/0.776 21.99/0.811 25.81/0.846 via multi-task collaboration
@@
$\text{TUR}_{\text{Transweather}}$ [131] AAAI’25 38M 29.68/0.966 33.09/0.952 30.40/0.869 26.63/0.815 23.02/0.838 28.56/0.888 a novel loss function
AdaIR [79] ICLR’25 29M 30.53/0.978 38.02/0.981 31.35/0.889 28.12/0.858 23.00/0.845 30.20/0.910 frequency mining and modulation
DA-RCOT [233] TPAMI’25 50M 30.96/0.975 37.87/0.980 31.23/0.888 28.68/0.872 23.25/0.836 30.40/0.911 as an optimal transport problem
ABAIR [234] arXiv’24 41M 33.46/0.983 38.18/0.983 31.38/0.898 29.00/0.878 24.20/0.865 31.24/0.921 a three-phase approach
MoCE-IR [153] CVPR’25 25M 30.48/0.974 38.04/0.982 31.34/0.887 30.05/0.899 23.00/0.852 30.58/0.919 mixture-of-complexity-experts framework
$\text{DCPT}_{\text{PromptIR}}$ [236] ICLR’25 35M 30.72/0.977 37.32/0.978 31.32/0.885 28.84/0.877 23.35/0.840 30.31/0.911 learn to classify degradation
RamIR [237] APPL INTELL’25 22M 31.09/0.9787 37.56/0.979 31.44/0.886 28.82/0.878 22.02/0.828 30.18/0.910 prompt-driven Mamba-based
D3Net [244] arXiv’25 38M 32.57/0.965 38.35/0.973 31.73/0.860 32.70/0.851 26.49/0.857 32.36/0.901 cross-domain and dynamic decomposition
Cat-AIR [203] arXiv’25 - 30.88/0.978 38.21/0.982 31.38/0.890 28.91/0.876 23.46/0.848 30.57/0.915 content and task-aware framework
DSwinIR [238] arXiv’25 24M 30.09/0.975 37.77/0.982 31.34/0.885 29.17/0.879 22.64/0.843 30.19/0.913 deformable sliding window transformer
$\text{CPL}_{\text{PromptIR}}$ [116] arXiv’25 36M 30.82/0.978 38.20/0.983 31.41/0.887 28.72/0.874 23.65/0.855 30.55/0.915 contrastive prompt regularization
AnyIR [239] arXiv’25 6M 31.41/0.980 38.51/0.983 31.37/0.891 28.85/0.883 23.11/0.856 30.65/0.919 local-global gated mechanism
MIRAGE [240] arXiv’25 10M 31.45/0.980 38.92/0.985 31.41/0.892 28.10/0.858 23.59/0.858 30.68/0.914 within the SPD manifold space
BaryIR [241] arXiv’25 - 31.12/0.976 38.05/0.981 31.43/0.891 29.30/0.888 23.38/0.852 30.66/0.918 in continuous barycenter space
[…]
Perceive-IR [84] $\bigstar$ TIP’25 42M+86M 28.19/0.964 37.25/0.977 31.44/0.887 29.46/0.886 22.88/0.833 29.84/0.909 quality-aware degradation
VLU-Net [149] $\bigstar$ CVPR’25 35M+88M 30.84/0.980 38.54/0.982 31.43/0.891 27.46/0.840 22.29/0.833 30.11/0.905 vision-language gradient descent-driven
[…]
TABLE IV: Performance Comparisons of AiOIR Models on All-Weather Dataset [29]. Below the dividing line in the table is AiOIR methods.
@@
Histoformer [253] 32.16/0.926 32.08/0.939 33.06/0.944 32.43/0.936 17M
LoRA-IR [155] 32.28/0.930 32.62/0.945 33.39/0.949 32.76/0.941 -
[…]
MWFormer [206] 30.92/0.908 30.27/0.912 31.91/0.927 31.03/0.916 170M
$\text{TUR}_{\text{Transweather}}$ [131] 30.62/0.909 29.75/0.907 31.61/0.933 30.66/0.916 38M
[…]
DSwinIR [238] 32.58/0.931 32.76/0.950 32.88/0.947 32.74/0.943 24M
$\mathrm{DA}^{2}\mathrm{Diff}$ [254] 31.42/0.916 31.58/0.939 33.01/0.945 31.67/0.933 -
HOGformer [255] 32.41/0.930 32.89/0.946 32.72/0.945 32.67/0.940 17M
$\text{CPL}_{\text{PromptIR}}$ [116] 32.27/0.928 32.16/0.942 32.73/0.943 32.39/0.938 36M
MODEM [256] 32.52/0.929 33.10/0.941 33.01/0.943 32.87/0.938 20M
For Setting3 in Tab. IV, MODEM [256] achieves the strongest overall performance by the Morton-order 2D-selective-scan module, with an average PSNR/SSIM of 32.87/0.938. It is closely followed by DSwinIR [238] and LoRA-IR [155], which deliver averages of 32.74/0.943 and 32.76/0.941, respectively, demonstrating the effectiveness of deformable window attention and mixture-of-experts. By contrast, early methods such as All-in-One [53] yield a much lower average PSNR of 28.05 dB, highlighting the substantial gains achieved by these newer, specialized architectures.
For Setting4 in Tab. V, although Transweather [29] carries the larger parameter count, it struggles to deliver competitive performance. In contrast, though more parameter-efficient, AirNet still falls short in overall restoration quality in the WeatherStream dataset. TKMANet [63] and WGWS-Net [65] show a notable improvement in both metrics, balancing effectiveness and efficiency. Zhu et al. [65] first construct the Real-World Dataset for image restoration under multiple weather conditions of real scenes. Specifically, we test on SPA+ [187] for deraining, RealSnow [65] for desnowing, and REVIDE [183] for dehazing. DSwinIR achieves superior performance on this dataset. Overall, these results demonstrate a clear progression in multi-weather restoration, suggesting that the field is advancing towards more sophisticated and robust techniques across diverse conditions.
For Setting5 in Tab. VI, we categorize existing methods into four settings: (a) One-to-One trains on isolated degradations; (b) One-to-Many trains on multiple isolated degradations and tests on them jointly; (c) One-to-Composite trains and tests on both isolated and composite degradations; and (d) Step-by-Step trains on isolated degradations and tests on their combinations step by step. Among the One-to-Composite methods, OneRestore [31], which introduces a novel dataset CDD-11 encompassing both single and composite degradations, achieves a good balance between accuracy and efficiency. In comparison, the AllRestorer model [257] achieves the best performance by adaptive weights and composite scene descriptors. For the Step-by-Step, $\text{CoR}_{\text{OneRestore}}$ demonstrates the advantage of chain of restoration, even only requiring training on $n$ isolated degradations instead of the full $12^{n}-$ 1 combinations. Overall, models in the one-to-composite and step-by-step categories show a certain degree of generalization and robustness, making them worthy of further exploration.
TABLE V: Performance Comparisons of AiOIR Models on Real-World Weather-Related Datasets.
@@
the Real-World Dataset [65]
TKMANet [63] 20.10/0.85 37.32/0.97 29.37/0.88 28.93/0.90 29M
TransWeather [29] 17.33/0.82 33.64/0.93 26.92/0.86 26.71/0.86 38M
$\text{TUR}_{\text{Transweather}}$ [131] 20.38/0.88 39.78/0.98 29.72/0.91 29.96/0.92 38M
WGWS-Net [65] 29.46/0.85 38.94/0.98 33.61/0.93 34.01/0.92 6M
DSwinIR [238] 30.14/0.89 40.60/0.98 33.80/0.93 34.85/0.93 24M
TABLE VI: Performance Comparison on CDD-11 Dataset. Methods are categorized into One-to-One, One-to-Many, and One-to-Composite.
  Type Method Venue & Year PSNR $\uparrow$ SSIM $\uparrow$ Params
  One-to-One MPRNet [225] CVPR’21 25.47 0.856 16M
MIRNetv2 [249] TPAMI’22 25.37 0.854 6M
Restormer [50] CVPR’22 26.99 0.865 26M
DGUNet [250] CVPR’22 25.33 0.844 18M
NAFNet [52] ECCV’22 26.22 0.796 17M
SRUDC [259] ICCV’23 27.64 0.860 7M
Fourmer [260] ICML’23 23.44 0.789 6M
OKNet [261] AAAI’24 26.33 0.861 6M
One-to-Many AirNet [54] CVPR’22 23.75 0.814 9M
TransWeather [29] CVPR’22 23.13 0.804 38M
WeatherDiff [64] TPAMI’23 24.09 0.799 83M
PromptIR [58] NeurIPS’23 25.90 0.850 36M
WGWS-Net [65] CVPR’23 26.96 0.861 6M
HAIR [205] arXiv’24 27.85 0.866 29M
One-to-Composite OneRestore [31] ECCV’24 28.47 0.878 6M
[…]
MoCE-IR-S [153] CVPR’25 29.05 0.881 11M
$\text{DCPT}_{\text{NAFNet}}$ [236] ICLR’25 28.84 0.891 68M
VL-UR [262] $\bigstar$ arXiv’25 28.76 0.879 -
MIRAGE [240] arXiv’25 29.33 0.887 10M
Step-by-Step $\text{CoR}_{\text{HAIR}}$ [161] arXiv’24 28.33 0.869 30M
$\text{CoR}_{\text{OneRestore}}$ [161] arXiv’24 28.84 0.879 8M